find the derivative(t+1)2 b) h(t) = 7735 (3 + 5t)5
Answer #1

sol: Given . 7). htt) = (t+1) ** t + st) The derivative of hlt) is nlt) = at dere (test) I = Aply the Quotient sale (1+09445- Peso (Cs) (+*+st) - Anly the power sule & sum oule - #19) (ECC) +s4 kt) = 5w4 (84 451) = sut (3+75) substitute back uz tis